# Quadratic dispersion relations in gapless frustration-free systems

### v1

@@ III.3 Degeneracy in gapped cases
The discussion above can be readily extended to the cubic lattice in three dimensions. The variational energy of the plane-wave state $|\Phi_{(k_{x},k_{y},k_{z})}\rangle\coloneqq L^{-3/2}\sum_{x,y,z=1}^{L}e^{i(k_{x}x+k_{y}y+k_{z}z)}\hat{s}_{(x,y,z)}^{-}|{\Phi_{0}}\rangle$ is simply given by $E(k_{x})+E(k_{y})+E(k_{z})$.
Finally, let us discuss whether this analysis can be generalized to other form of lattices. For example, the nearest-neighbor interaction on the triangular lattice contains a diagonal interaction
$\sum_{x,y=1}^{L}\hat{Q}_{(x,y),(x+1,y-1)}.$ (57)
More generally, when an interaction $\hat{Q}_{(x,y),(x+d_{x},y+d_{y})}$ among the spins at $(x,y)$ and $(x+d_{x},y+d_{y})$ is added, the variational energy obtains a term
$E(k_{x}d_{x}+k_{y}d_{y})$
$=\sum_{x,y=1}^{L}\langle\Phi_{(k_{x},k_{y})}|\hat{Q}_{(x,y),(x+d_{x},y+d_{y})}|\Phi_{(k_{x},k_{y})}\rangle.$ (58)
@@
Hence, the variational state generically becomes gapped unless $E(k_{0}(d_{x}+d_{y}))=0$. When $k_{0}=0$ (for example, in the ferromagnetic Heisenberg model), this condition can be easily satisfied, but otherwise some fine-tuning is required.
V 1D examples with next nearest-neighbor interactions
In Sec. III, we showed that frustration-free Hamiltonians for $s=1/2$ spin chains with nearest-neighbor interactions have a common ground state and a low-energy variational state. This observation was the key in demonstrating the softness of the dispersion relation in higher dimensional models in Sec. IV.
@@
Table 1: Comparison of gapless frustration-free Hamiltonians for $s=1/2$ chain. $|{\Phi_{0}}\rangle$ [(15)] is the fully polarized state and $|{\Psi_{k}}\rangle$ [Eq. (16)] is the plane-wave state of single spin flip.
@@
XYZ MG model: (78), (82) with $\theta\neq 0$ 2 $O\big{(}k^{2})$
In this section, we discuss $s=1/2$ spin chains with longer-range interactions. We assume that the Hamiltonian has the following form
$\hat{H}^{\mathrm{pbc}}\coloneqq\sum_{x=1}^{L}\hat{Q}_{x,x+1,x+2},\quad\hat{H}^{\mathrm{obc}}\coloneqq\sum_{x=1}^{L-2}\hat{Q}_{x,x+1,x+2}.$ (59)
Here, $\hat{Q}_{x,x+1,x+2}$ is a projector acting on the three spins at $x$, $x+1$, and $x+2$. The interaction range of the Hamiltonian is 2. Unlike the only nearest-neighbor interaction case in the previous section, our discussion here is based on examples and hence is not comprehensive. However, these examples clearly demonstrate difficulties in the direct extension of our results to more general cases. Namely, the fully polarized state $|{\Phi_{0}}\rangle$ may not be a ground state and the plane-wave state $|\Psi_{k}\rangle$ may not be a low-energy state. We instead construct a low-energy variational state which does not take the form $\sum_{x}e^{ikx}\hat{O}_{x}|\Phi_{0}\rangle$. We summarize our results in Table 1.
V.1 W state
Let us start with the uncle Hamiltonian constructed by the fully polarized state and the W state $|W\rangle\coloneqq(|100\rangle+|010\rangle+|001\rangle)/\sqrt{3}$ Fernández-González et al. (2015). The kernel of the Hamiltonian is
$\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}=\mathrm{span}\big{\{}|000\rangle,|W\rangle\big{\}}.$ (60)
The spin representation of the Hamiltonian is found to be
$\hat{Q}_{x,x+1,x+2}$ $\coloneqq\frac{3}{4}\hat{\mathbbm{1}}-\frac{\hat{s}_{x}^{z}+\hat{s}_{x+1}^{z}+\hat{s}_{x+2}^{z}}{3}+2\hat{s}_{x}^{z}\hat{s}_{x+1}^{z}\hat{s}_{x+2}^{z}$
$\quad-\frac{2}{3}\Big{(}\frac{1}{2}\hat{\mathbbm{1}}+\hat{s}_{x}^{z}\Big{)}\hat{\bm{s}}_{x+1}\cdot\hat{\bm{s}}_{x+2}$
$\quad-\frac{2}{3}\Big{(}\frac{1}{2}\hat{\mathbbm{1}}+\hat{s}_{x+1}^{z}\Big{)}\hat{\bm{s}}_{x+2}\cdot\hat{\bm{s}}_{x}$
$\quad-\frac{2}{3}\Big{(}\frac{1}{2}\hat{\mathbbm{1}}+\hat{s}_{x+2}^{z}\Big{)}\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+1}.$ (61)
The fully-polarized state $|{\Phi_{0}}\rangle$ and the zero momentum state of a single spin flip $|\Psi_{0}\rangle$ are the two ground states of this model regardless of the boundary condition. The state $|{\Psi_{k}}\rangle$ in Eq. (16) is an exact eigenstate of $\hat{H}^{\mathrm{pbc}}$ as expected by the U(1) symmetry. The eigen energy is given by
$E(k)$ $=\big{|}\hat{Q}_{1,2,3}(|100\rangle+e^{ik}|010\rangle+e^{2ik}|001\rangle)\big{|}^{2}$
$=\big{|}\hat{Q}_{1,2,3}((e^{ik}-1)|010\rangle+(e^{2ik}-1)|001\rangle)\big{|}^{2}$
$=2(1-\cos k)\big{|}\hat{Q}_{1,2,3}(|010\rangle+(e^{ik}+1)|001\rangle)\big{|}^{2}$
$=\frac{8}{3}(2+\cos k)\sin^{2}(k/2).$ (62)
As a related model, let us discuss the bulk Hamiltonian for the Fredkin spin chain Salberger and Korepin (2017):
$\hat{Q}_{x,x+1,x+2}^{\prime}$ $\coloneqq\Big{(}\frac{1}{2}\hat{\mathbbm{1}}+\hat{s}_{x}^{z}\Big{)}\Big{(}\frac{1}{4}\hat{\mathbbm{1}}-\hat{\bm{s}}_{x+1}\cdot\hat{\bm{s}}_{x+2}\Big{)}$
$\quad+\Big{(}\frac{1}{4}\hat{\mathbbm{1}}-\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+1}\Big{)}\Big{(}\frac{1}{2}\hat{\mathbbm{1}}-\hat{s}_{x+2}^{z}\Big{)}.$ (63)
The kernel of this Hamiltonian is
$\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}^{\prime}$ $=\mathrm{span}\Big{\{}|000\rangle,|100\rangle,|001\rangle+|010\rangle,$
[…]
Since $\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}^{\prime}\supset\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}$, the min-max principle implies that the Fredkin spin chain is gapless with dispersion equal to or softer than the one implied by Eq. (62). Indeed, $|{\Psi_{k}}\rangle$ is an exact eigenstate with eigenenergy
$E^{\prime}(k)$ $=1-\cos k=2\sin^{2}(k/2).$ (65)
In these two examples, the fully polarized state $|{\Phi_{0}}\rangle$ is one of the ground states and the plane-wave state $|\Psi_{k}\rangle$ is a low-energy state with the dispersion $E(k)=O(k^{2})$.
V.2 Greenberger–Horne–Zeilinger uncle Hamiltonian
@@
Next, we discuss the uncle Hamiltonian for the Greenberger–Horne–Zeilinger (GHZ) state Fernández-González et al. (2015). The kernel of the Hamiltonian is
$\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}$
$=\mathrm{span}\Big{\{}|000\rangle,|111\rangle,|100\rangle+|110\rangle,|001\rangle+|011\rangle\Big{\}}.$ (66)
The spin representation the Hamiltonian is
$\hat{Q}_{x,x+1,x+2}$ $\coloneqq\frac{1}{2}\hat{\mathbbm{1}}-\hat{s}_{x+1}^{z}(\hat{s}_{x}^{z}+\hat{s}_{x+2}^{z})$
$\quad-2\hat{s}_{x+1}^{x}\Big{(}\frac{1}{4}-\hat{s}_{x}^{z}\hat{s}_{x+2}^{z}\Big{)}.$ (67)
@@
To write down ground states and low-energy excitations of this model, let us define
$|n\rangle\coloneqq|\underbrace{0\cdots 0}_{L-n}\underbrace{1\cdots 1}_{n}\rangle,\quad|\bar{n}\rangle\coloneqq|\underbrace{1\cdots 1}_{L-n}\underbrace{0\cdots 0}_{n}\rangle.$ (68)
for $n=1,2,\cdots,L-1$. For example, $|n\rangle$ satisfies
$\hat{H}^{\mathrm{pbc}}|n\rangle-2|n\rangle$
$=-\frac{1}{2}(|n+1\rangle+|n-1\rangle+\hat{T}|n+1\rangle+\hat{T}^{-1}|n-1\rangle)$ (69)
@@
$\hat{H}^{\mathrm{pbc}}|n\rangle-2|n\rangle$
=$=\begin{cases}-\frac{1}{2}(|2\rangle+\hat{T}|2\rangle)&(n=1),\\ -\frac{1}{2}(|L-2\rangle+\hat{T}^{-1}|L-2\rangle)&(n=L-1)\end{cases}$ (70)
for $n=1$ and $L-1$. Based on these expressions, the ground states of $\hat{H}^{\mathrm{obc}}$ are found to be
@@
The latter two are not consistent with the PBC and are hence not ground states of $\hat{H}^{\mathrm{pbc}}$.
Although the fully-polarized state $|{\Phi_{0}}\rangle$ is one of the ground states, the simple variational state $|{\Psi_{k}}\rangle$ is not a low-energy state:
$\langle\Psi_{k}|\hat{H}^{\mathrm{pbc}}|{\Psi_{k}}\rangle=2.$ (72)
We instead consider a plane-wave state of two domain walls:
$|\Psi_{k,\ell}\rangle\coloneqq\frac{\sqrt{2}}{L}\sum_{n=1}^{L-1}\sum_{m=0}^{L-1}e^{ikm}e^{-ikn/2}\sin(\tfrac{\pi n\ell}{L})\hat{T}^{m}|n\rangle$ (73)
for $\ell=1,2,\cdots,L-1$. Using $\hat{T}|\Psi_{k,\ell}\rangle=e^{-ik}|\Psi_{k,\ell}\rangle$ and $[\hat{T},\hat{H}^{\mathrm{pbc}}]=0$, one can readily show that $|\Psi_{k,\ell}\rangle$ is an exact eigenstate of $\hat{H}^{\mathrm{pbc}}$ with the eigenvalue
$E_{\ell}(k)$ $=2[1-\cos(k/2)\cos(\tfrac{\pi\ell}{L})]$
$=4\sin^{2}(\tfrac{\pi\ell}{2L})+4\cos(\tfrac{\pi\ell}{L})\sin^{2}(k/4).$ (74)
The first term in the last expression is $O(L^{-2})$, which vanishes in the thermodynamic limit. The second term is $O(k^{2})$.
Despite the presence of gapless excitations, in this model, the correlation function of the form
@@
decays exponentially with the system size for any choice of local operators $\hat{\mathcal{O}}_{\bm{x}}$ and $\hat{\mathcal{O}}^{\prime}_{\bm{y}}$. Here $\hat{G}$ is the projector onto the ground states of $\hat{H}^{\mathrm{pbc}}$. This is because both of the two ground states of $\hat{H}^{\mathrm{pbc}}$ (the first two states in Eq. (71)) are product states.
V.3 XYZ MG model
As our last example, let us discuss variations of the MG model. The local Hamiltonian of the original model is given by
$\hat{Q}_{x,x+1,x+2}$
$\coloneqq\frac{1}{2}\hat{\mathbbm{1}}+\frac{2}{3}(\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+1}+\hat{\bm{s}}_{x+1}\cdot\hat{\bm{s}}_{x+2}+\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+2}).$ (76)
Any singlet state $|s\rangle_{x,x+1}\coloneqq(|01\rangle-|10\rangle)/\sqrt{2}$ belongs to the kernel of this Hamiltonian.
$\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}$
$=\mathrm{span}\Big{\{}|010\rangle-|001\rangle,|100\rangle-|010\rangle,$
@@ V.3 XYZ MG model
$\hat{Q}_{x,x+1,x+2}^{\prime}$
$\coloneqq J_{0}\hat{\mathbbm{1}}+\sum_{a=x,y,z}J_{a}(\hat{s}_{x}^{a}\hat{s}_{x+1}^{a}+\hat{s}_{x+1}^{a}\hat{s}_{x+2}^{a}+\hat{s}_{x}^{a}\hat{s}_{x+2}^{a})$ (78)
@@
the model exhibits additional ground state degeneracy. Following Ref. Palle and Benton (2021), we set $J_{0}=\frac{1}{4}$ and
@@
Unless $\theta=0$, the fully polarized state $|{\Phi_{0}}\rangle$ is not a ground state of the system. The $\theta=\pm\pi/4$ cases are unitary equivalent to the model discussed in Ref. Saito and Hotta (2024). The plane-wave state $|{\Psi_{k}}\rangle$ in Eq. (16) is not a low-energy state unless $\theta=0$; indeed, we find
$\langle\Psi_{k}|\hat{H}^{\mathrm{pbc}}|{\Psi_{k}}\rangle=\frac{\cos^{2}\theta}{3}(1+2\cos k)^{2}+(L-3)\sin^{2}\theta.$ (86)
@@
$|\tilde{\Psi}_{k}\rangle=\frac{1}{\sqrt{Ln_{k}}}\sum_{m=0}^{L-1}e^{ikm}\hat{T}^{m}|D\rangle,$ (87)
@@
$\langle\tilde{\Psi}_{k}|\hat{H}|\tilde{\Psi}_{k}\rangle$ $=\frac{\big{\|}\hat{Q}_{2,3,4}(|D\rangle+\hat{T}^{2}|D\rangle)\big{\|}^{2}}{n_{k}}=\frac{2\sin^{2}k}{3n_{k}}.$ (89)
This is $O(k^{2})$, which is consistent with the result of Ref. Palle and Benton (2021) on Kagome lattice. Hence, our conjecture still holds in this example as well, despite the fact that $|{\Phi_{0}}\rangle$ is not a ground state and $|{\Psi_{k}}\rangle$ is not gapless.

### v2

@@ III.2 Rank 2 case
which satisfies (4). This discussion can be readily extended to the cubic lattice. The variational energy of the plane-wave state $|\Psi_{\bm{k}}\rangle\coloneqq L^{-3/2}\sum_{x,y,z=1}^{L}e^{i\bm{k}\cdot\bm{r}}\hat{s}_{(x,y,z)}^{-}|{\Phi_{0}}\rangle$ is simply given by $E(k_{x})+E(k_{y})+E(k_{z})$.
@@
Finally, let us investigate possible generalization to other form of lattices. For example, the nearest-neighbor interaction on the triangular lattice contains a diagonal interaction $\sum_{x,y=1}^{L}\hat{Q}_{(x,y),(x+1,y-1)}$. More generally, when an interaction $\hat{Q}_{(x,y),(x+d_{x},y+d_{y})}$ among the spins at $(x,y)$ and $(x+d_{x},y+d_{y})$ is added, the variational energy obtains a term $E(k_{x}d_{x}+k_{y}d_{y})$. Hence, the plane-wave state generically becomes gapped unless $E(k_{0}(d_{x}+d_{y}))=0$. When $k_{0}=0$ (e.g., in the ferromagnetic Heisenberg model), this condition can be easily satisfied, but otherwise some fine-tuning is required.
V Examples
In this section we discuss several examples for which our general discussing in Sec. III and IV are not applicable. We summarize our results in Table 1.
@@ V Examples
Table 1: Comparison of gapless FF Hamiltonians for $s=1/2$ chain. $|{\Phi_{0}}\rangle=|00\cdots\rangle$ is the fully polarized state and $|{\Psi_{k}}\rangle$ [Eq. (6)] is the plane-wave state of single spin flip. The XYZ MG model is discussed in Sec. IV of SM.
@@
XYZ MG model with $\theta\neq 0$: (29) 2 $O\big{(}k^{2})$
V.1 Kinetic Ising model
Let us consider an $s=1/2$ spin model defined on a $d$-dimensional cubic lattice: $\hat{H}=\sum_{\bm{r}}\hat{H}_{\bm{r}}$, where
$\hat{H}_{\bm{r}}=\frac{1}{2\cosh(J\sum_{\bm{r}^{\prime}\in B_{\bm{r}^{\prime}}}\hat{\sigma}_{\bm{r}^{\prime}}^{z})}\Big{(}e^{-J\hat{\sigma}_{\bm{r}}^{z}{}\sum_{\bm{r}^{\prime}\in B_{\bm{r}}}\hat{\sigma}_{\bm{r}^{\prime}}^{z}}-\hat{\sigma}_{\bm{r}}^{x}\Big{)}.$ (18)
Here, $\hat{\sigma}_{\bm{r}}^{a}$ ($a=x,y,z$) are the Pauli matrices related to spin operators by $\hat{s}_{\bm{r}}^{a}=(1/2)\hat{\sigma}_{\bm{r}}^{a}$ and $B_{\bm{r}}$ is the nearest neighbors of $\bm{r}$. This model is associated with the classical Ising model with the Boltzmann weight $\propto e^{J\sum_{\langle\bm{r},\bm{r}^{\prime}\rangle}\sigma_{\bm{r}}^{z}\sigma_{\bm{r}^{\prime}}^{z}}$ Masaoka et al. (2024). (The temperature $T$ is included in the definition of $J$.) The $\mathbb{Z}_{2}$ symmetry $\prod_{\bm{r}}\hat{\sigma}_{\bm{r}}^{x}$ of the model is spontaneously broken only in the limit $J\to\infty$, which corresponds to the zero temperature limit, in one dimension and for $J$ above a critical value $J_{c}$ in higher dimensions.
V.1.1 1D
In one dimension, $\hat{H}_{x}$ is a projector $\hat{Q}_{x-1,x,x+1}$ acting on the three spins at $x-1$, $x$, and $x+1$, which can be expressed as
$\hat{Q}_{x-1,x,x+1}$
$=\frac{1}{2}\hat{\mathbbm{1}}-c_{2}\hat{s}_{x}^{z}(\hat{s}_{x-1}^{z}+\hat{s}_{x+1}^{z})-2\hat{s}_{x}^{x}\Big{(}\frac{c_{1}}{4}-c_{3}\hat{s}_{x-1}^{z}\hat{s}_{x+1}^{z}\Big{)}$ (19)
with $c_{1}=\frac{2\cosh^{2}J}{\cosh(2J)}$, $c_{2}=\tanh(2J)$, and $c_{3}=\frac{2\sinh^{2}J}{\cosh(2J)}$.
@@ V.1 Kinetic Ising model
In the $J\to\infty$ limit, all three parameters $c_{1}$, $c_{2}$, $c_{3}$ become $1$ and the model reduces to the uncle Hamiltonian for the Greenberger–Horne–Zeilinger (GHZ) state Fernández-González et al. (2015). The kernel of $\hat{Q}_{x,x+1,x+2}$ is given by
$\mathrm{span}\Big{\{}|000\rangle,|111\rangle,|100\rangle+|110\rangle,|001\rangle+|011\rangle\Big{\}}.$ (20)
Figure 2: (a) The all spin-up state $|\Phi_{0}\rangle$ of 1D kinetic Ising model. (b) An excited state $|\Psi_{k,\ell}\rangle$ with two domain walls in Eq. (24). (c) A ground state of the 2D kinetic Ising model with straight domain walls. (d) One-dimensional gapless excitations on the surface of domain walls that corresponds to (b).
@@
To write down ground states and low-energy excitations of this model, let us introduce $|\bar{n}\rangle\coloneqq|\underbrace{1\cdots 1}_{L-n}\underbrace{0\cdots 0}_{n}\rangle$ in addition to $|n\rangle\coloneqq|\underbrace{0\cdots 0}_{L-n}\underbrace{1\cdots 1}_{n}\rangle$. For example, $|n\rangle$ satisfies
$\hat{H}|n\rangle-2|n\rangle$
$=-\frac{1}{2}(|n+1\rangle+|n-1\rangle+\hat{T}_{1}|n+1\rangle+\hat{T}_{1}^{-1}|n-1\rangle)$ (21)
@@
$\hat{H}|n\rangle-2|n\rangle$
=$=\begin{cases}-\frac{1}{2}(|2\rangle+\hat{T}_{1}|2\rangle)&(n=1),\\ -\frac{1}{2}(|L-2\rangle+\hat{T}_{1}^{-1}|L-2\rangle)&(n=L-1)\end{cases}$ (22)
for $n=1$ and $L-1$. Based on these expressions, the ground states under OBC are found to be
@@
The latter two are not consistent with the PBC and are hence not ground states of $\hat{H}$.
Although the fully-polarized state $|{\Phi_{0}}\rangle$ [Fig 2(a)] is one of the ground states, the plane-wave state in Eq. (6) is not a low-energy state because $\langle\Psi_{k}|\hat{H}|{\Psi_{k}}\rangle=2$. We instead consider a plane-wave state of two domain walls illustrated in Fig 2(b):
$|\Psi_{k,\ell}\rangle\coloneqq\frac{\sqrt{2}}{L}\sum_{n=1}^{L-1}\sum_{m=0}^{L-1}e^{ikm}e^{-ikn/2}\sin(\tfrac{\pi n\ell}{L})\hat{T}_{1}^{m}|n\rangle$ (24)
for $\ell=1,2,\cdots,L-1$. We find that $|\Psi_{k,\ell}\rangle$ is an exact eigenstate of $\hat{H}$ and $\hat{T}_{1}$ with the energy eigenvalue
$E_{\ell}(k)=4\sin^{2}\Big{(}\frac{\pi\ell}{2L}\Big{)}+4\cos\Big{(}\frac{\pi\ell}{L}\Big{)}\sin^{2}\Big{(}\frac{k}{4}\Big{)}.$ (25)
The first term is $O(L^{-2})$, which vanishes in the thermodynamic limit. The second term is $O(k^{2})$.
Despite the presence of gapless excitations, all the correlation functions of the form
@@
decay exponentially with the system size for any choice of local operators $\hat{\mathcal{O}}_{\bm{x}}$ and $\hat{\mathcal{O}}^{\prime}_{\bm{y}}$. Here $\hat{G}$ is the projector onto the ground states of $\hat{H}$. This is because both of the two ground states of $\hat{H}$ (the first two states in Eq. (23)) are product states.
V.1.2 2D
In two dimensions, the critical point $J_{c}$ is known to be $(1/2)\log(1+\sqrt{2})=0.440687\cdots$. Above $J>J_{c}$, the $\mathbb{Z}_{2}$ symmetry is spontaneously broken and many quasi-degenerate ground states appear, whose finite size splitting decay exponentially with the system size. Hence one may think that this model is a counterexample to our first conjecture that states the absence of finite size splitting in gapped FF system. However, it is not the case because the model is known to be gapless above $J>J_{c}$ Miyashita and Takano (1985); Henkel and Pleimling (2010). For example, in the limit $J\to\infty$, not only all spin up state and all spin down states, there are many other ground states with straight domain walls as illustrated in Fig 2(c)  [^2] The width of each domain must be greater than one. The one-dimensional excitation on the surface of a domain wall [Fig 2(d)] that precisely takes the form $|\Psi_{k,\ell}\rangle$ in Eq. (24) is a gapless excitation with the same excitation energy $E_{\ell}(k)$ in Eq. (25).
V.2 XYZ MG model
As another nontrivial example, let us discuss a variation of the MG model Sriram Shastry and Sutherland (1981); Shastry and Sutherland (1981); Gerhardt et al. (1998); Xu et al. (2021); Palle and Benton (2021). The local Hamiltonian of the original model is given by
$\hat{Q}_{x,x+1,x+2}\coloneqq\frac{1}{2}\hat{\mathbbm{1}}+\frac{2}{3}(\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+1}+\hat{\bm{s}}_{x+1}\cdot\hat{\bm{s}}_{x+2}+\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+2}).$ (27)
Any singlet state $|s\rangle_{x,x+1}\coloneqq(|01\rangle-|10\rangle)/\sqrt{2}$ belongs to the kernel of $\hat{Q}_{x,x+1,x+2}$:
$\mathrm{span}\Big{\{}|010\rangle-|001\rangle,|100\rangle-|010\rangle,$
@@ V.2 XYZ MG model
$\hat{Q}_{x,x+1,x+2}^{\prime}\coloneqq\frac{1}{4}\hat{\mathbbm{1}}+\sum_{a=x,y,z}J_{a}(\hat{s}_{x}^{a}\hat{s}_{x+1}^{a}+\hat{s}_{x+1}^{a}\hat{s}_{x+2}^{a}+\hat{s}_{x}^{a}\hat{s}_{x+2}^{a})$ (29)
@@
the model exhibits additional ground state degeneracy. Following Ref. Palle and Benton, 2021, we set
@@
Unless $\theta=0$, the fully polarized state $|{\Phi_{0}}\rangle$ is not a ground state of the system. The $\theta=\pm\pi/4$ cases are unitary equivalent to the model discussed in Ref. Saito and Hotta, 2024. The plane-wave state $|{\Psi_{k}}\rangle$ in Eq. (4) of the main text is not a low-energy state unless $\theta=0$; indeed, we find
$\langle\Psi_{k}|\hat{H}|{\Psi_{k}}\rangle=\frac{\cos^{2}\theta}{3}(1+2\cos k)^{2}+(L-3)\sin^{2}\theta.$ (37)
@@
$|\tilde{\Psi}_{k}\rangle=\frac{1}{\sqrt{Ln_{k}}}\sum_{m=0}^{L-1}e^{ikm}\hat{T}_{1}^{m}|D\rangle,$ (38)
@@
$\langle\tilde{\Psi}_{k}|\hat{H}|\tilde{\Psi}_{k}\rangle$ $=\frac{\big{\|}\hat{Q}_{2,3,4}(|D\rangle+\hat{T}_{1}^{2}|D\rangle)\big{\|}^{2}}{n_{k}}=\frac{2\sin^{2}k}{3n_{k}}.$ (40)
This is $O(k^{2})$, which is consistent with the result of Ref. Palle and Benton, 2021 on Kagome lattice. Hence, our conjecture still holds in this example as well, despite the fact that $|{\Phi_{0}}\rangle$ is not a ground state and $|{\Psi_{k}}\rangle$ is not gapless.